OpenStudy (anonymous):

Please help! EASY Algebra Use Linear combination method for the following questions. 1. 3x + y = 4 2x + y = 4 A. -3, 12 B. -1, 7 C. 0, 4 D. -3, 13 2. 3x + 2y = -7 5x - 4y = -19 A. no solution B. -3,1 C. infinitive number of solutions D. -4, 2 Determine how many solutions each equation has. 1. 3x - 9y = 0 -x + 3y = -4 A. 1 B. 2 C. infinitive D. none 2. 6x - 9y = -9 2x - 3y = -6 Solve. 1. 2x - y = 5 3y + 27 = 4 A. -1, -7 B. 2, -1 C.0, -2 D.5,5 2. x + y = 3 2x + 2y = 6 A. 1.5, 1.5 B.none C.infinitive D.0,3

OpenStudy (anonymous):

3x + 2y = -7 5x - 4y = -19 A. none B. -3, 1 C. infinitive D. -4, 2

iYuko (iyuko):

just graph it.

iYuko (iyuko):

Go to Math On demand

OpenStudy (anonymous):

I figured out the first two... 1. c 2. b

pooja195 (pooja195):

Ok so far so good Can you do the steps to #3 show me what u know

OpenStudy (anonymous):

ok just a sec

OpenStudy (anonymous):

3x - 9y = 0 -x + 3y = -4